Determine the slope-intercept form of a linear equation, given the listed attributes: a) Slope = -2 and y-intercept = (0,10) b) Slope = -3 and (4, -2) lies on line c) Slope = 0 and (2,4) lies on line d) (3, -2) and (-12,1) lies on line e) (20, 240) and (15,450) lies on line


1
Expert's answer
2021-08-02T16:51:42-0400

a) If slope is "-2" and y-intercept is "(0,10)" then the slope-intercept form of a linear equation is "y=-2x+10."


b) If slope "-3" and "(4, -2)" lies on line then the slope-intercept form of a linear equation is "y=-3x+b," where "-2=-3\cdot 4+b," that is "b=10." Therefore, "y=-3x+10."


c) If slope is and "(2,4)" lies on line then the slope-intercept form of a linear equation is "y=4."


d) The points (3, -2) and (-12,1) lies on line given by the following equation:


"\frac{x-3}{-12-3}=\frac{y-(-2)}{1-(-2)}," which is equivalent to the following equations


"\frac{x-3}{-15}=\frac{y+2}{3}"


"y+2=\frac{x-3}{-5}"


"y=-2-\frac{x}{5}+\frac{3}{5}"


Therefore, the slope-intercept form of a linear equation is


"y=-\frac{1}{5}x-\frac{7}{5}."


e) The points "(20, 240)" and "(15,450)" lies on the line given by the following equation:


"\frac{x-15}{20-15}=\frac{y-450}{240-450}," which is equivalent to the following equations


"\frac{x-15}{5}=\frac{y-450}{-210}"


"y-450=-42(x-15)"


Therefore, the slope-intercept form of a linear equation is


"y=-42x+1080"